OpenStudy (anonymous):

2x(x+2)=3 Solve using the quadratic formula

OpenStudy (08surya):

break it

OpenStudy (08surya):

and then apply sridharacharya formula

Nnesha (nnesha):

first distribute x+2 by 2x

Nnesha (nnesha):

do you understand?? do you know how distribute this?? @steven_g

OpenStudy (solomonzelman):

people are making their lives more complicated. \(\normalsize\color{black}{ 2x(x+2)=3\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ 2x^2+4x=3\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ x^2+2x=1.5\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ x^2+2x+1=2.5\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{black}{ (x+1)^2=2.5\LARGE\color{white}{ \rm │ }}\)